OpenStudy (anonymous):

Solve: 8x2 = -24x. Your answer should include the complete solution set. {0, 3} {-3, 3} {-3, 0} {-3, 0, 3}

OpenStudy (anonymous):

I NEED HELP

hero (hero):

Hint: add 24x to both sides

OpenStudy (anonymous):

First put the equation in general form 8x^2+24x=0 Then take out the common factor 8x 8x(x+3)=0 Using the Zero Factor Theorem Either 8x=0 or x+3=0 Hence x=0 or x=-3

OpenStudy (anonymous):

Is this a multiple choice question?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Well what do you think the answer is?

OpenStudy (anonymous):

-3,0

OpenStudy (anonymous):

AM I RIGHT???

OpenStudy (anonymous):

Yes...:). I like to draw it on the number line, if this helps with visualizing.

OpenStudy (anonymous):

It starts at -3 and ends at 0.

OpenStudy (anonymous):

OKAY

OpenStudy (anonymous):

IM ABOUT TO PUT ANOTHER QUESTION UP

OpenStudy (anonymous):

Okay, don't forget to select best answer to the person who you think deserves it most :).

OpenStudy (anonymous):

And try working the question first and then let us know where you get stuck or if you have no clue then let us know so we can guide you and give you pointers :). All the best!!!